external work equals to 4 x 10^(-7) J is needed to move a -5-nC charge at a constant speed to a point at which the potential is -20V. what is the potential at the initial point?
1
Expert's answer
2012-10-30T10:19:06-0400
A=(U2-U1)q
U1=U2-A/q
U1= -20V- 4 x 10^(-7) J/(-5-nC) = -20V-80V= -100V